### v1

@@ Abstract
This paper addresses distributed optimization problems in non-i.i.d. scenarios, focusing on the interplay between communication and computation efficiency. To this end, we propose FlexGT, a flexible snapshot gradient tracking method with tunable numbers of local updates and neighboring communications in each round. Leveraging a unified convergence analysis framework, we prove that FlexGT achieves a linear or sublinear convergence rate depending on objective-specific properties–from (strongly) convex to nonconvex–and the above-mentioned tunable parameters. FlexGT is provably robust to the heterogeneity across nodes and attains the best-known communication and computation complexity among existing results. Moreover, we introduce an accelerated gossip-based variant, termed Acc-FlexGT, and show that with prior knowledge of the graph, it achieves a Pareto-optimal trade-off between communication and computation. Particularly, Acc-FlexGT achieves the optimal iteration complexity of $\tilde{\mathcal{O}}\left(L/\epsilon+L\sigma^{2}/\left(n\epsilon^{2}\sqrt{1-\sqrt{\rho_{W}}}\right)\right)$ for the nonconvex case, matching the existing lower bound up to a logarithmic factor, and improves the existing results for the strongly convex case by a factor of $\tilde{\mathcal{O}}\left(1/\sqrt{\epsilon}\right)$, where $\epsilon$ is the targeted accuracy, $n$ the number of nodes, $L$ the Lipschitz constant, $\rho_{W}$ the spectrum gap of the graph, and $\sigma$ the stochastic gradient variance. Numerical examples are provided to demonstrate the effectiveness of the proposed methods.
keywords:
Distributed optimization, data heterogeneity, complexity analysis, gradient tracking, accelerated communication, Pareto-optimal
{MSCcodes}
90C06, 90C25, 90C26
@@ 1 Introduction
(1) $\underset{x\in\mathbb{R}^{p}}{\min}f\left(x\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{:=f_{i}\left(x\right)}{\underbrace{\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[f_{i}\left(x;\xi_{i}\right)\right]}}},$
@@ 1.1 Literature Review
Decentralized stochastic gradient descent (DSGD) [39] is a simple yet efficient algorithm for solving problem (1). It performs stochastic gradient descent locally while communicating with neighbors to achieve consensus, which enables data parallelism and avoids communication bottlenecks in centralized architectures [32]. However, the native DSGD methods often struggle in scenarios with non-independent and identically distributed (non-i.i.d.) local datasets [24, 12]. For (1), such a scenario corresponds to that the distribution $\mathcal{D}_{i}$ is different. This data heterogeneity induces gradient drift in the sense that local gradients computed at each node deviate from the global gradient direction, necessitating decaying stepsizes to ensure exact convergence [35, 24]. This requirement limits the applicability of DSGD in many real-world tasks, as non-i.i.d. data distributions are common.
@@
Many efforts have been made to address this issue of data heterogeneity in distributed optimization. Gradient tracking (GT) methods have gained popularity for mitigating the effect [51, 11, 41, 23, 46, 48]. Notably, Xu et al. [51] proposed AugDGM, which introduces an auxiliary variable at each node to track the average of the local gradients, enabling exact convergence with a constant stepsize. This method has been extended to broader settings, including time-varying graphs [34] and directed graphs [38, 40], with convergence guarantees. Pu et al. [37] extended GT to the stochastic setting, introducing distributed stochastic gradient tracking (DSGT). They proved that DSGT achieves linear convergence to a neighborhood of the optimal solution, determined by the stochastic gradient noise. This extension enhances the applicability of GT, for instance, to many machine learning tasks. Although GT methods are robust to data heterogeneity, they exhibit a stronger dependence on the graph compared to DSGD [13, 18], leading to increased communication overhead. The consensus alternating direction method of multipliers (C-ADMM) [27] and its inexact variant [6] also exhibit robustness to data heterogeneity. However, they require either solving a subproblem exactly at each iteration, which incurs high computation overhead, or assuming a well-structured objective function to enable linear approximations, which limits their applicability. These observations highlight the need for more efficient algorithms, especially in resource-constrained scenarios that involve both communication and computation.
@@
To mitigate the high communication burden, many improved GT-based methods have been developed [36, 30, 43, 10, 42]. A popular strategy is to skip some communication steps and perform local updates only. For instance, FedAvg [32] (also known as Local SGD [44]) is widely used in federated learning [8, 17], where periodic communication occurs between server and worker nodes, with partial node participation to reduce communication overhead. In decentralized settings, Nguyen et al. [36] introduced LU-GT, a variant of DSGT that uses multiple local updates, and provided communication complexity results matching those of Local SGD for nonconvex objectives. Subsequently, Liu et al. [30] proposed the gradient-sum-tracking method K-GT, which reduces communication complexity by achieving lower stochastic gradient variance. However, both [36] and [30] overlook the side effect of how the local updates increase computation complexity due to worse inter-node consensus. Another approach is to randomly or periodically communicate over a subset of the graph’s edges to reduce communication [39, 25, 26, 52]. These methods can be seen as approximate asynchronous update schemes, which often require specific connectivity assumptions [39] or particular topologies, such as exponential graphs [52].
@@
In contrast to directly skipping certain communications as in Local SGD [44], another strategy involves using accelerated consensus protocols. For instance, Chebyshev acceleration [43, 10] and accelerated gossip [28, 31, 54] methods enable the distributed algorithm to behave more like a centralized algorithm. In particular, Song et al. [43] introduced an optimal gradient tracking (OGT) method with Chebyshev acceleration for deterministic optimization problems with smooth and strongly convex objective functions. Leveraging prior knowledge of the network topology, they demonstrated that a limited number of communication steps per round could enable the distributed algorithm to replicate centralized updates. Building on this, the stochastic variant SS-DSGT [10] was proposed, achieving lower communication complexity compared to DSGT. For nonconvex cases, Lu et al. [31] developed an accelerated DSGT method called DeTAG, achieving optimal iteration complexity. However, the effects of these acceleration techniques on both communication and computation remain unclear. Although communication quantization and compression can be used to further reduce the amount of data transmitted in each round [20, 42], their impacts on algorithmic accuracy must be carefully considered. For a comprehensive overview of other efforts to improve communication efficiency, readers are referred to recent survey papers [47, 5].
@@
Some recent efforts have focused on both the communication and computation efficiency of distributed optimization algorithms. Berahas et al. [3] proposed a distributed gradient descent method with multiple communication steps in each round, named NEAR-DGD, and evaluated the performance of the algorithm using a weighted combination of the number of communication and computation steps. In [2], the authors explored the use of multiple local updates and inter-node communication in each round with a deterministic GT method, showing that increasing the number of communication steps improves the convergence rate in strongly convex settings. However, in distributed stochastic optimization, balancing communication and computation is more challenging. This is because stochastic gradient noise significantly influences the convergence rate, making the results from deterministic methods [3, 2] inapplicable to stochastic settings. Although the results in [15, 29] highlight the impact of communication and computation protocols on iteration complexity, they do not provide a specific trade-off strategy between these two factors and are all affected by the presence of data heterogeneity. Additionally, most of the studies focus on specific types of objective functions, such as strongly convex or nonconvex objectives, or rely on strict assumptions, such as bounded gradients [29]. We summarize and compare the most relevant algorithms in Table 1.
@@
Ref. Objectives Comm. Comp. Stoch. Acc. Non-i.i.d.
DSGT[37, 18] SC, Co, NC Single Single
@@
LU-GT[36] NC Single Multi.
K-GT [30] NC Single Multi.
DeTAG [31] NC Multi. Single
GTA [2] SC Multi. Multi.
S-NEAR-DGD[15] SC Multi. Multi.
DFL [29] NC Multi. Multi.
(Acc-)FlexGT (ours) SC, Co, NC Multi. Multi.
  • Note: SC, Co, and NC denote strongly convex, convex, and nonconvex objective functions, respectively. “Single” and “Multi.” indicate that the algorithm uses single or multiple local communication (Comm.) or computation (Comp.) steps per round, respectively. “Stoch.” refers to the use of stochastic gradients. “Acc.” refers to communication acceleration.
@@ 1.2 Contributions
This paper solves the distributed stochastic optimization problem (1) in scenarios with non-i.i.d local datasets. We introduce the FlexGT algorithm and illustrate its trade-offs between communication and computation. Pareto-optimal trade-offs are further achieved through an accelerated variant, termed Acc-FlexGT. The main contributions are summarized as follows:
  • We propose FlexGT, a flexible gradient tracking method that allows adjustable numbers of local updates and communications in each round, enabling trade-offs between communication and computation. By introducing a so-called snapshot point within the GT step, FlexGT mitigates the inconsistency between GT variables during local updates, resulting in improved convergence. Moreover, by integrating an accelerated gossip communication protocol, we develop Acc-FlexGT, which significantly enhances both communication and computation efficiency, and, to the best of our knowledge, is the first to achieve Pareto-optimal trade-offs between these two aspects, under the assumption of prior knowledge of the graph.
  • We develop a unified convergence analysis framework for FlexGT and Acc-FlexGT, covering objective functions that are strongly convex, convex, or nonconvex. We establish linear or sublinear convergence rates with the best-known communication and computation complexity in most settings, in terms of their dependencies on problem-specific parameters, compared to existing results (Theorem 1 and Corollary 1). In particular, we prove that Acc-FlexGT achieves the optimal iteration complexity, matching the existing lower bound in the nonconvex case up to a logarithmic factor, and improves existing results in the strongly convex case (Corollary 2). To the best of our knowledge, this work provides the first complexity analysis of an accelerated GT method in a convex setting. This unified analysis framework can be extended to a broader class of distributed optimization algorithms, which is of independent interest.
@@ 1.3 Paper Organization
The remainder of this paper is structured as follows. Section 2 formulates the distributed stochastic optimization problem over networks under standard assumptions and introduces the FlexGT and Acc-FlexGT algorithms. Section 3 then presents the main convergence result of the algorithms for strongly convex, convex, and nonconvex objective functions, showing the dependency on both the amount of communication and computation. Section 4 provides the proof of the result. Numerical experiments on synthetic and real-world datasets are given in Section 5 to validate the effectiveness of FlexGT and Acc-FlexGT. Finally, Section 6 concludes the paper.
@@ 2 Problem Formulation and Algorithm Design
(2) $\underset{\mathbf{x}\in\mathbb{R}^{n\times p}}{\min}F\left(\mathbf{x}\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{:=f_{i}\left(x_{i}\right)}{\underbrace{\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[f_{i}\left(x_{i};\xi_{i}\right)\right]}},}\quad\mathrm{s.t.}\,\,x_{i}=x_{j},\quad\forall i,j\in\left[n\right]$
where $\mathbf{x}:=\left[x_{1},x_{2},\dots,x_{n}\right]\in\mathbb{R}^{n\times p}$ denotes the collection of the local copies of the decision variables from all nodes. The constraints guarantee that problems (1) and (2) are equivalent. Moreover, the agents are connected over a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Here, $\mathcal{V}=\{1,2,...,n\}$ represents the set of nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of edges consisting of ordered pairs $(i,j)$ representing the communication link from node $j$ to node $i$. For node $i\in[n]$, we define $\mathcal{N}_{i}=\{j|\left(i,j\right)\in\mathcal{E}\}$ as the set of its neighboring nodes (including $i$ itself).
2.1 Assumptions
@@
Particularly, we say that $f_{i}$ is $\mu$-strongly convex if $\mu>0$; otherwise, we refer to it as convex.
Assumption 2 (Smoothness).
@@
In our algorithms, we assume that each agent $i$ obtains an unbiased noisy gradient of the form $\nabla f_{i}\left(x;\xi_{i}\right)$ by querying a stochastic oracle ($\mathcal{S}\mathcal{O}$) with $\xi_{i}\sim\mathcal{D}_{i}$ at each iteration, where ${D}_{i}$ is an arbitrary distribution.
Assumption 3 (Bounded variance).
@@
Assumption 4 (Graph connectivity).

### v2

@@ Abstract
This paper addresses distributed stochastic optimization problems under non-i.i.d. data, focusing on the inherent trade-offs between communication and computational efficiency. To this end, we propose FlexGT, a flexible snapshot gradient tracking method that enables tunable numbers of local updates and neighbor communications per round, thereby adapting efficiently to diverse system resource conditions. Leveraging a unified convergence analysis framework, we derive tight communication and computational complexity for FlexGT with explicit dependence on objective properties and certain tunable parameters. Moreover, we introduce an accelerated variant, termed Acc-FlexGT, and prove that, with prior knowledge of the graph, it achieves Pareto-optimal trade-offs between communication and computation. Particularly, in the nonconvex case, Acc-FlexGT achieves the optimal iteration complexity of $\tilde{\mathcal{O}}\left(\left(L\sigma^{2}\right)/\left(n\epsilon^{2}\right)+L/\left(\epsilon\sqrt{1-\sqrt{\rho_{W}}}\right)\right)$ and optimal communication complexity of $\tilde{\mathcal{O}}\left(L/\left(\epsilon\sqrt{1-\sqrt{\rho_{W}}}\right)\right)$ for appropriately chosen numbers of local updates, matching existing lower bounds up to logarithmic factors. And, it improves the existing results for the strongly convex case by a factor of $\tilde{\mathcal{O}}\left(1/\sqrt{\epsilon}\right)$, where $\epsilon$ is the targeted accuracy, $n$ the number of nodes, $L$ the Lipschitz constant, $\rho_{W}$ the connectivity of the graph, and $\sigma$ the stochastic gradient variance. Numerical experiments corroborate the theoretical results and demonstrate the effectiveness of the proposed methods.
{IEEEkeywords}
Distributed optimization, data heterogeneity, accelerated communication, complexity analysis
@@ 1 Introduction
$\underset{x\in\mathbb{R}^{p}}{\min}f\left(x\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{:=f_{i}\left(x\right)}{\underbrace{\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[f_{i}\left(x;\xi_{i}\right)\right]}}},$ (1)
@@ 1.1 Literature Review
Decentralized stochastic gradient descent (DSGD) [38] is a simple yet efficient algorithm for solving Problem (1). It performs stochastic gradient descent locally while communicating with neighbors to achieve consensus, which enables data parallelism and avoids communication bottlenecks in centralized architectures [31]. However, the native DSGD methods often struggle in scenarios with non-independent and identically distributed (non-i.i.d.) local datasets [23, 12]. For Problem (1), such a scenario corresponds to the case where the local data distributions $\mathcal{D}_{i}$ differ across agents. This data heterogeneity induces gradient drift in the sense that local gradients computed at each node deviate from the global gradient direction, necessitating decaying stepsizes to ensure exact convergence [34, 23]. This requirement limits the applicability of DSGD in many real-world tasks, as non-i.i.d. data distributions are common.
@@
Many efforts have been made to address the issue of data heterogeneity in distributed optimization. Gradient tracking (GT) methods have gained popularity for mitigating the effect [51, 11, 40, 22, 45, 47]. Notably, Xu et al. [51] proposed AugDGM, which introduces an auxiliary variable at each node to track the average of the local gradients, enabling exact convergence with a constant stepsize. This method has been extended to broader settings, including time-varying graphs [33] and directed graphs [37, 39], with convergence guarantees. Pu et al. [36] extended GT to the stochastic setting, introducing distributed stochastic gradient tracking (DSGT). They proved that DSGT achieves linear convergence to a neighborhood of the optimal solution, determined by the stochastic gradient noise. This extension enhances the applicability of GT, for instance, to many machine learning tasks. Although GT methods are robust to data heterogeneity, they exhibit a stronger dependence on the graph compared to DSGD [13, 19], leading to increased communication overhead. The consensus alternating direction method of multipliers (C-ADMM) [26] and its inexact variant [6] also exhibit robustness to data heterogeneity. However, they require either solving a subproblem exactly at each iteration, which incurs high computation overhead, or assuming a well-structured objective function to enable linear approximations, which limits their applicability. These observations highlight the need for more efficient algorithms, especially in resource-constrained scenarios that involve both communication and computation.
@@
To mitigate the high communication burden, many improved GT-based methods have been developed [35, 41, 42, 10, 49, 29]. A popular strategy is to skip some communication steps and perform local updates only. For instance, FedAvg [31] (also known as Local SGD [43]) is widely used in federated learning [8, 17], where periodic communication occurs between server and worker nodes, with partial node participation to reduce communication overhead. In decentralized settings, Nguyen et al. [35] introduced LU-GT, a variant of GT that uses multiple local updates, and provided communication complexity results matching those of Local SGD for nonconvex objectives. The effectiveness of multiple local updates under mild data heterogeneity is further demonstrated in [49]. Subsequently, Liu et al. [29] proposed the gradient-sum-tracking method K-GT, which reduces communication complexity by achieving lower stochastic gradient variance. However, both [35] and [29] mainly focus on communication efficiency and do not explicitly consider how increasing the number of local updates may raise computational complexity due to weakened inter-node consensus. Another approach is to randomly or periodically communicate over a subset of the graph’s edges to mitigate the impact of single-point communication bottleneck and stragglers [38, 24, 25, 52]. These methods can be seen as approximate asynchronous update schemes, which often require specific connectivity assumptions [38] or particular topologies, such as exponential graphs [52].
@@
In contrast to directly skipping certain communications, as in Local SGD [43], another strategy involves using accelerated consensus protocols. For instance, Chebyshev acceleration [42, 10] and accelerated gossip [27, 30, 54] methods enable the distributed algorithm to behave more like a centralized algorithm. In particular, Song et al. [42] introduced an optimal gradient tracking (OGT) method with Chebyshev acceleration for deterministic optimization problems with smooth and strongly convex objective functions. Leveraging prior knowledge of the network topology, they demonstrated that a limited number of communication steps per round could enable the distributed algorithm to replicate centralized updates. Building on this, the stochastic variant SS-DSGT [10] was proposed, achieving lower communication complexity compared to DSGT. For nonconvex cases, Lu et al. [30] developed an accelerated DSGT method called DeTAG, achieving optimal iteration complexity with large batches of stochastic gradients. Although communication quantization and compression can be used to further reduce the amount of data transmitted in each round [18, 41], their impacts on algorithmic accuracy must be carefully considered. For a comprehensive overview of other efforts to improve communication efficiency, readers are referred to the recent survey papers [46, 5].
@@
Some recent efforts have focused on both the communication and computational efficiency of distributed optimization algorithms. Berahas et al. [3] proposed a distributed gradient descent method with multiple communication steps in each round, named NEAR-DGD, and evaluated the performance of the algorithm using a weighted combination of the number of communication and computation steps. In [2], the authors explored the use of multiple local updates and inter-node communication in each round with a deterministic GT method, showing that increasing the number of communication steps improves the convergence rate for the strongly convex objective functions. However, in distributed stochastic optimization, balancing communication and computation is more challenging, as stochastic gradient noise critically affects the convergence rate and renders results from deterministic methods inapplicable [3, 2]. Although the results in [15, 28] highlight the impact of communication and computation protocols on iteration complexity, the trade-off between these two factors is not explicitly characterized, particularly in the presence of data heterogeneity. Additionally, most of the studies focus on specific types of objective functions, such as strongly convex or nonconvex objectives, or rely on strict assumptions, such as bounded gradients [28]. We summarize and compare the most relevant algorithms in Table 1.
@@
Reference Objectives Communication Computation Stochastic Acceleration Non-i.i.d.
DSGT[36, 19] SC, Cvx, NC Single Single
@@
LU-GT[35] NC Single Multiple
K-GT [29] NC Single Multiple
DeTAG [30] NC Multiple Single
GTA [2] SC Multiple Multiple
S-NEAR-DGD[15] SC Multiple Multiple
DFL [28] NC Multiple Multiple
(Acc-)FlexGT (ours) SC, Cvx, NC Multiple Multiple
  • Note: SC, Cvx, and NC denote strongly convex, convex, and nonconvex objective functions, respectively. “Single” and “Multiple” indicate that the algorithm uses single or multiple local communication or computation steps per round, respectively. “Stochastic” refers to the use of stochastic gradients. “Acceleration” refers to communication acceleration.
@@ 1.2 Contributions
This paper solves the distributed stochastic optimization problem (1) in scenarios with non-i.i.d local datasets. We introduce the FlexGT algorithm and illustrate the inherent trade-offs between communication and computation. Pareto-optimal trade-offs (cf., Definition 1) are further achieved through an accelerated variant, termed Acc-FlexGT. The contributions are summarized as follows:
  • We propose FlexGT, a flexible gradient tracking method that allows adjustable numbers of local updates and communications in each round, enabling better adaptation to heterogeneous communication and computation resources in distributed systems. By introducing a snapshot point within the GT step, FlexGT mitigates the inconsistency between GT variables during local updates, leading to improved convergence. Moreover, by integrating an accelerated gossip communication protocol, we develop Acc-FlexGT, which significantly enhances both communication and computational efficiency when the graph connectivity parameter is known a priori. To the best of our knowledge, Acc-FlexGT is the first algorithm to achieve Pareto-optimal trade-offs between communication and computation from a multi-objective optimization perspective, jointly minimizing these two costs under given resource budgets.
  • We develop a unified convergence analysis framework for FlexGT and Acc-FlexGT, covering objective functions that are strongly convex, convex, or nonconvex. By employing refined error decomposition and decoupling techniques, we establish linear or sublinear convergence rates with the best-known communication and computational complexity in most settings, explicitly characterizing the dependence on problem-specific parameters, compared to existing GT algorithms (Theorem 1 and Corollary 1). In particular, we show that Acc-FlexGT achieves the optimal iteration complexity and optimal communication complexity in the nonconvex case with appropriately chosen numbers of local updates, matching existing lower bounds up to logarithmic factors (Corollary 2). Moreover, Acc-FlexGT improves the best-known results in the strongly convex case by a factor of $\tilde{\mathcal{O}}(1/\sqrt{\epsilon})$. This unified analysis framework can be extended to a broader class of algorithms, which is of independent interest.
@@ 1.3 Paper Organization
The remainder of this paper is structured as follows. Section 2 formulates the distributed stochastic optimization problem over networks under standard assumptions and introduces the FlexGT and Acc-FlexGT algorithms. Section 3 presents the main convergence result of the algorithms for strongly convex, convex, and nonconvex objective functions, showing the dependency on both the amount of communication and computation. Section 4 provides the proof of the result. Numerical experiments on synthetic and real-world datasets are given in Section 5 to validate the effectiveness of FlexGT and Acc-FlexGT. Finally, Section 6 concludes the paper.
@@ 2 Problem Formulation and Algorithm Design
$\underset{\mathbf{x}\in\mathbb{R}^{n\times p}}{\min}$ $F\left(\mathbf{x}\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{:=f_{i}\left(x_{i}\right)}{\underbrace{\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[f_{i}\left(x_{i};\xi_{i}\right)\right]}},}$ (2)
$\mathrm{s.t.}\,\,x_{i}=x_{j},\quad\forall i,j\in\left[n\right]$
where $\mathbf{x}:=\left[x_{1},x_{2},\dots,x_{n}\right]^{\top}\in\mathbb{R}^{n\times p}$ denotes the collection of the local copies of the decision variables from all nodes. The constraints guarantee that problems (1) and (2) are equivalent. Moreover, the agents are connected over a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Here, $\mathcal{V}=\{1,2,...,n\}$ represents the set of nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of edges consisting of ordered pairs $(i,j)$ representing the communication link from node $j$ to node $i$. For node $i\in[n]$, we define $\mathcal{N}_{i}=\{j|\left(i,j\right)\in\mathcal{E}\}$ as the set of its neighboring nodes (including $i$ itself).
Figure 1: Flow diagram of the computation and communication of FlexGT. The variable $\psi_{i,l}$, $i=1,\dots,n$, represents the collection of $x_{i,l}$, $y_{i,l}$, and $z_{i,l}$. Only $x_{i,l}$ and $y_{i,l}$ are communicated. The communication is weighted by the matrix $W$.
2.1 Assumptions and Definitions
@@
Particularly, we say that $f_{i}$ is $\mu$-strongly convex if $\mu>0$; otherwise, we refer to it as being convex.
Assumption 2 (Smoothness)
@@
Moreover, the overall objective function is bounded from below, i.e., $f\left(x\right)\geqslant f^{*}:=\mathrm{inf}_{x\in\mathbb{R}^{p}}\,\,f\left(x\right)>$-$\infty$.
In our algorithms, we assume that each agent $i$ obtains an unbiased noisy gradient of the form $\nabla f_{i}\left(x;\xi_{i}\right)$ by querying a stochastic oracle ($\mathcal{S}\mathcal{O}$) with $\xi_{i}\sim\mathcal{D}_{i}$ at each iteration, where $\mathcal{D}_{i}$ is the local data distribution. This is typically achieved by uniformly random sampling on each node.
Assumption 3 (Bounded variance)
@@
Assumption 4 (Graph connectivity)
@@
In this work, we aim to jointly minimize communication and computational complexities, measured respectively by the total number of inter-node communications and local stochastic gradient evaluations needed to achieve a prescribed accuracy. This naturally leads to a multi-objective optimization problem that admits a set of Pareto-optimal solutions, as defined below.
Definition 1 (Pareto-optimality [4, 7])
For a multi-objective optimization problem that minimizes $\left\{g_{j}\left(\omega\right)\right\}_{j\in\left[m\right]}$, a point $\omega$ is Pareto-optimal if there does not exist another point $\omega^{\prime}$ such that: $\forall j\in[m]$, $g_{j}\left(\omega^{\prime}\right)\leqslant g_{j}\left(\omega\right)$; and $\exists j\in\left[m\right]$, $g_{j}\left(\omega^{\prime}\right)<g_{j}\left(\omega\right)$. The set of Pareto-optimal points is called the Pareto frontier.
In what follows, we refer to the Pareto-optimal points that jointly optimize communication and computational complexity in terms of tunable parameters as Pareto-optimal trade-offs.
It should be noted that the above multi-objective optimization perspective is fundamentally different from the existing iteration-complexity–based analyses [30, 54], which characterizes the minimum number of algorithmic iterations required to reach a target accuracy, under a given function class, algorithm class, and stochastic oracle model. Such settings do not explicitly account for communication and computational costs per iteration, both of which are, indeed, influenced by the use of multiple communication rounds, multiple local updates, or large-batch stochastic gradients.